# Off-Path TCP Hijacking in Wi-Fi Networks: A Packet-Size Side Channel Attack

### v1

@@ Abstract
In this paper, we unveil a fundamental side channel in Wi-Fi networks, specifically the observable frame size, which can be exploited by attackers to conduct TCP hijacking attacks. Despite the various security mechanisms (e.g., WEP and WPA2/WPA3) implemented to safeguard Wi-Fi networks, our study reveals that an off path attacker can still extract sufficient information from the frame size side channel to hijack the victim’s TCP connection. Our side channel attack is based on two significant findings: (i) response packets (e.g., ACK and RST) generated by TCP receivers vary in size, and (ii) the encrypted frames containing these response packets have consistent and distinguishable sizes. By observing the size of the victim’s encrypted frames, the attacker can detect and hijack the victim’s TCP connections. We validate the effectiveness of this side channel attack through two case studies, i.e., SSH DoS and web traffic manipulation. Furthermore, we conduct extensive measurements to evaluate the impact of our attack on real-world Wi-Fi networks. We test 30 popular wireless routers from 9 well-known vendors, and none of these routers can protect victims from our attack. Also, we implement our attack in 80 real-world Wi-Fi networks and successfully hijack the victim’s TCP connections in 69 (86%) evaluated Wi-Fi networks. We have responsibly disclosed the vulnerability to the Wi-Fi Alliance and proposed several mitigation strategies to address this issue.
@@ 1 Introduction
Nowadays, public Wi-Fi networks are widely available in various places, such as airports, coffee shops, hotels, and libraries. To prevent a malicious supplicant (i.e., an attacker) from eavesdropping in wireless channels, security measures (e.g., WEP, WPA2 and WPA3) have been proposed to encrypt Wi-Fi networks at the link layer [^1] Our study of 80 real-world Wi-Fi networks reveals that 64 of them are encryption enabled.. As a result, even if an attacker successfully sniffs wireless frames in the shared Wi-Fi channel, it cannot get useful information (e.g., the random sequence and acknowledgment numbers of TCP connections) except for the 802.11 frame header.
However, in this paper, we demonstrate that the encrypted frame size constitutes a reliable side channel that can be exploited by attackers to conduct TCP hijacking attacks. Precisely, we discover that TCP packets can be identified by analyzing the size of the encrypted wireless frames, thus allowing an attacker residing in the same Wi-Fi network to infer the state of a victim TCP connection. By exploiting this side channel (i.e., the encrypted frame size), the attacker can infer the random sequence and acknowledgment numbers of the victim TCP connection. Consequently, the attacker can pretend to be one peer of the victim connection to either terminate the connection or inject malicious data into the connection, i.e., hijacking the connection completely.
Our attack consists of four steps. First, the attacker accesses a public Wi-Fi network and probes alive supplicants in the WLAN. The attacker crafts ARP requests in the WLAN to identify alive supplicants. By collecting the ARP replies, the attacker can obtain the <MAC, IP> address pair of each alive supplicant which is also a potential victim client of our TCP hijacking attack. Then through analyzing the MAC address field of the captured wireless frames in the shared Wi-Fi channels, the attacker can filter the encrypted frames belonging to the victim client. If the Wi-Fi network provides multiple access channels, the attacker can scan all Wi-Fi channels to filter the victim’s frames. Once the victim’s frames are sniffed, the attacker gains a potent side channel to conduct the TCP hijacking attack.
@@
We conduct a comprehensive measurement study to show that our attack can be performed to cause serious damage in the real world, e.g., terminating a victim SSH connection or poisoning a web traffic within 335 seconds. We test 30 popular wireless routers from 9 well-known vendors, and we discover that none of these routers can protect victims from our attack. Besides, we evaluate our attack in 80 real-world Wi-Fi networks, including most popular Wi-Fi scenarios (e.g., Wi-Fi networks in coffee shops, bookstores, shopping malls, and restaurants). The experimental results show that 69 (86%) out of the 80 evaluated Wi-Fi networks are vulnerable to our TCP hijacking attack.
@@ 2.1 Frame Format and Security Mechanisms in Wi-Fi Network
The 802.11 Frame Layout. Figure 1 shows the layout of an 802.11 frame. Firstly, the Frame Control (FC) field contains several flags and defines the type of frame. The Type and Subtype fields together identify the function of the frame. There are currently three types (i.e., management, control, and data frames) and more than 50 subtypes defined in 802.11 specifications. In our attack, the attacker needs to monitor the victim’s TCP packets which will be encapsulated into 802.11 frames with type 2 and subtype 8 in Wi-Fi networks. To identify the victim’s encrypted frames, the attacker needs to analyze the addresses of the 802.11 frames. There are four address fields in the MAC frame format. These fields are used to indicate the basic service set identifier (BSSID), source address (SA), destination address (DA), transmitting address (TA), and receiving address (RA). Certain frames might not contain some of the address fields. Certain address field usage is specified by the relative position of the address field (1–4) within the MAC header, independent of the type of address present in that field. Specifically, the Address 1 field always identifies the intended receiver(s) of the frame, and the Address 2 field, where present, always identifies the transmitter of the frame [3]. In our attack, the attacker can identify the victim supplicant’s encrypted frames through addresses 1 (RA) and address 2 (TA). After filtering the victim’s encrypted frames, the attacker needs to further analyze the payload size of the encrypted frames. The payload (i.e., MSDU in Figure 1) of a normal data frame contains the upper layer data (e.g., TCP packets). The MSDU typically starts with an LLC/SNAP header and is protected by cryptographic encapsulation mechanisms (i.e., TKIP, CCMP, and GCMP). In this paper, we refer to the encrypted frame size as the MSDU size.
Figure 1: Layout of the 802.11 frame.
@@ 2.2 Challenge ACK Mechanism in TCP
Our attack leverages the trigger conditions of the challenge ACK mechanism in two distinct ways. Firstly, when a receiver detects an incoming SYN packet within an established TCP connection, regardless of the sequence number, it responds by sending an ACK (referred to as the challenge ACK) to the remote peer. This ACK serves as a challenge for the remote peer to confirm the loss of the previous connection and the initiation of a new connection. Only the legitimate peer will receive this ACK and respond with a RST segment containing the correct sequence number, derived from the ACK field of the challenge ACK packet, in the event of connection loss. Consequently, a spoofed SYN packet will generate an additional ACK, which will be disregarded by the peer as a duplicate ACK and will have no impact on the established connection. We will demonstrate how this challenging condition can be exploited to detect a victim TCP connection.
Secondly, the receiver employs a verification process for the acknowledgment number of each TCP segment to prevent blind data injection attacks. Acceptance of an acknowledgment number for any data segment is contingent upon its falling within the range of ($SND.UNA-SND.WND,SND.NXT$), where $SND.UNA$ represents the sequence number of the first unacknowledged octet, $SND.WND$ denotes the maximum window size observed by the receiver from the sender, and $SND.NXT$ is the next sequence number to be sent. If the acknowledgment number of the segment ($SEG.ACK$) is in the range ($SND.UNA-(2^{31}-1),SND.UNA-SND.WND$), the receiver responds with a challenge ACK. If the $SEG.ACK$ is greater than $SND.NXT$, the receiver silently discards this TCP segment. That can be exploited by attackers to infer the acceptable acknowledgment number, as described in Sec 4.3.3.
@@ 2.3 TCP Options
Selective Acknowledgment Option. The TCP selective acknowledgment (SACK) is specified in RFC 2018 [23]. It is an optional feature that is typically enabled by default in the majority of TCP implementations. The SACK option is particularly recommended for networks experiencing frequent packet loss or packet reordering. Its utilization can significantly improve the performance and reliability of TCP connections in such environments. When a receiver detects a TCP segment with a sequence number that has already been acknowledged as outdated, it responds by sending a SACK-ACK to notify the sender. The SACK option will enable the sender to selectively retransmit lost packets based on the information provided by the receiver. However, if the sequence number of the received TCP segment has not yet been acknowledged, the receiver will reply with an ACK packet, which may have a different frame size (as shown in Table 1) or may not respond at all, depending on the acknowledgment number of the segment. In this paper, we will show that attackers in Wi-Fi networks can differentiate between these situations and thus infer the sequence number of a target TCP connection by analyzing the size of sniffed wireless frames.
@@
3 Threat Model
Refer to caption (a) Network scanning and detecting TCP connections. Refer to caption (b) Inferring sequence and acknowledgment numbers.
@@ 3 Threat Model
Figure 2 illustrates the threat model of our off-path TCP hijacking attack in Wi-Fi networks. The AP encrypts the network traffic of its supplicants via security mechanisms, e.g., WPA2 or WPA3. A victim supplicant, such as a laptop or a smartphone, connects to the AP and establishes TCP connections with remote servers. The attacker, functioning as a regular supplicant without AP management privileges, utilizes multiple wireless network interface cards (WNICs). One (managed model) of these WNICs connects to the AP, while the others (monitor model) are utilized to sniff encrypted frames transmitted over the shared Wi-Fi channels. We make the assumption that the attacker has prior access to the target Wi-Fi network before performing our attack. This is a commonly accepted assumption in Wi-Fi hijacking scenarios, as highlighted in previous studies [22, 60].
@@ 4.1 Attack Overview
Step 1: Network Scanning. The attacker accesses a Wi-Fi network and scans the WLAN for potential victim supplicants. In this step, the attacker identifies the <MAC, IP> pair of the victim to monitor its encrypted frames.
@@
Step 3: Inferring Sequence Number. After detecting a victim TCP connection, the attacker sends forged TCP packets with guessed sequence numbers to the server. These manipulated TCP packets prompt the server to generate SACK-ACK responses, which will be sniffed by the attacker when they (i.e., 80-byte encrypted frames) are transmitted in the Wi-Fi channel. By monitoring the victim’s encrypted frames, the attacker can identify the correct sequence number of the target TCP connection.
@@
4.2 Network Scanning and Detecting TCP Connections
Network Scanning. The attacker first prepares the TCP hijacking attack from two aspects, i.e., obtaining the <MAC, IP> address pair of the victim and identifying the Wi-Fi channel used by the victim. The attacker actively sends ARP requests in the WLAN to detect other alive supplicants (i.e., the potential victim clients of our TCP hijacking attack). By observing the ARP responses, the attacker can learn the victim’s MAC address and IP address. With the victim’s MAC address, the attacker sniffs encrypted frames in the Wi-Fi channel and filters the victim’s frames based on address 1 (or address 2) in the 802.11 MAC header (see Figure 1). If the Wi-Fi network supports multiple accessed Wi-Fi channels, the attacker scans all Wi-Fi channels to identify the specific channel used by the victim. Subsequently, the attacker intercepts encrypted frames within the target Wi-Fi channel and filters out the victim’s frames.
Detecting TCP Connections. With intercepting and analyzing the victim’s encrypted frames, the attacker can identify the victim’s TCP connections. A TCP connection is recognized by four elements, i.e., [client IP address, client port number, server IP address, server port number]. Typically, the server IP address and server port number are publicly known to the attacker [58, 14]. The attacker needs to infer the client’s IP address and port number. In our attack, the client IP is obtained via ARP response. Thus, the only remaining element to determine is the client port number.
Given that a TCP connection was previously established by the legitimate user on a victim client using a source port p, the attacker impersonates as the client and sends forged SYN/ACK packets to the server. As per the challenge ACK mechanism described in RFC 5961, if the forged SYN/ACK packet contains the same client port number p, the server will respond with a challenge ACK to the client. This challenge ACK packet will be encapsulated into a 68-byte encrypted frame and sniffed by the attacker, during transmission from the AP to the client.
@@ 4.2 Network Scanning and Detecting TCP Connections
The attacker iterates through the above procedure by changing the client port number specified in the forged SYN/ACK packet. This procedure continues until the correct port number p is identified. Finally, the attacker identifies a target TCP connection operating on the four-tuple.
@@ 4.3 Inferring Sequence and Acknowledgment Numbers
In this section, we begin with a concise overview of the mechanism used to verify the sequence number and acknowledgment number of TCP segments. Next, we introduce a two-step approach to infer the exact sequence number and an acceptable acknowledgment number.
@@ 4.3.1 Verifying TCP Segment
According to RFC 9293 [18], upon receiving a TCP segment, the TCP receiver first performs a verification by comparing the sequence number ($SEG.SEQ$) specified in the TCP header with its receive window. In other words, the condition $RCV.NXT\leq SEG.SEQ\leq RCV.NXT+RCV.WND$ must be met, where $RCV.NXT$ denotes the next expected sequence number for an incoming segment, and $RCV.WND$ indicates the size of the receive window. Furthermore, as per the specification, the ACK flag is consistently set to true, except for the initial SYN packet used for connection establishment. If the ACK bit is disabled, the receiver will discard the segment. Therefore, when hijacking the target TCP connection, the attack must also infer an acceptable acknowledgment number, which cannot be evaded by deactivating the ACK flag.

### v4

@@ Abstract
In this paper, we unveil a fundamental side channel in Wi-Fi networks, specifically the observable frame size, which can be exploited by attackers to conduct TCP hijacking attacks. Despite the various security mechanisms (e.g., WEP and WPA2/WPA3) implemented to safeguard Wi-Fi networks, our study reveals that an off path attacker can still extract sufficient information from the frame size side channel to hijack the victim’s TCP connection. Our side channel attack is based on two significant findings: (i) response packets (e.g., ACK and RST) generated by TCP receivers vary in size, and (ii) the encrypted frames containing these response packets have consistent and distinguishable sizes. By observing the size of the victim’s encrypted frames, the attacker can detect and hijack the victim’s TCP connections. We validate the effectiveness of this side channel attack through two case studies, i.e., SSH DoS and web traffic manipulation. Precisely, our attack can terminate the victim’s SSH session in 19 seconds and inject malicious data into the victim’s web traffic within 28 seconds. Furthermore, we conduct extensive measurements to evaluate the impact of our attack on real-world Wi-Fi networks. We test 30 popular wireless routers from 9 well-known vendors, and none of these routers can protect victims from our attack. Besides, we implement our attack in 80 real-world Wi-Fi networks and successfully hijack the victim’s TCP connections in 75 (93.75%) evaluated Wi-Fi networks. We have responsibly disclosed the vulnerability to the Wi-Fi Alliance and proposed several mitigation strategies to address this issue.
@@ 1 Introduction
Nowadays, public Wi-Fi networks are widely available in various places, such as airports, coffee shops, hotels, and libraries. Serving as a prevalent method for Internet access, Wi-Fi networks have undergone substantial advancements in security mechanisms, progressing from WEP to WPA3, to counter various crypto-cracking attacks [40, 60, 64, 65]. Consequently, it becomes difficult for an off-path attacker to get useful information (e.g., the random sequence and acknowledgment numbers of TCP connections) from the encrypted Wi-Fi frames. Additionally, certain security policies (e.g., AP isolation and rogue AP detection [37, 33]) are proposed to counteract ARP poisoning and rogue APs. Moreover, recent efforts have rectified certain implementation vulnerabilities to thwart attackers from manipulating the router’s transmission queues [48], NAT mappings [69], and the next-hop routing via malicious ICMP redirects [23]. As a result, it poses a challenge for off-path attackers to hijack Wi-Fi network traffic.
However, in this paper, we demonstrate that the encrypted frame size constitutes a reliable side channel that can be exploited by attackers to conduct TCP hijacking attacks, even in Wi-Fi networks with AP isolation enabled. Precisely, we discover that TCP packets can be identified by analyzing the size of the encrypted wireless frames, thus allowing an attacker residing in the same Wi-Fi network to infer the state of the victim’s TCP connection. By exploiting this side channel (i.e., the encrypted frame size), the attacker can infer the random sequence and acknowledgment numbers of the victim’s TCP connection. Consequently, the attacker can pretend to be one peer of the victim’s connection to either terminate the connection or inject malicious data into the connection, i.e., hijacking the connection completely.
Our attack consists of four steps. First, the attacker accesses a public Wi-Fi network and probes alive supplicants in the WLAN. The attacker crafts ARP requests in the WLAN to identify alive supplicants [^1] Attackers can also identify alive supplicants by exploiting the DHCP mechanism, especially to circumvent the AP isolation mechanism enabled in Wi-Fi networks. Refer to Section 4.4 for more details.. By collecting the ARP replies, the attacker can obtain the $<MAC,IP>$ address pair of each alive supplicant which is also a potential victim client of our TCP hijacking attack. Then through analyzing the MAC address field of the captured wireless frames in the shared Wi-Fi channels, the attacker can filter the encrypted frames belonging to the victim client. If the Wi-Fi network provides multiple access channels, the attacker can scan all Wi-Fi channels to filter the victim’s frames. Once the victim’s frames are sniffed, the attacker gains a potent side channel to conduct the TCP hijacking attack.
@@
We conduct a comprehensive measurement study to show that our attack can be performed to cause serious damage in the real world, e.g., terminating a victim SSH connection or poisoning a web traffic within 28 seconds. We test 30 popular wireless routers from 9 well-known vendors, and we discover that none of these routers can protect victims from our attack. Besides, we evaluate our attack in 80 real-world Wi-Fi networks, including most popular Wi-Fi scenarios (e.g., Wi-Fi networks in coffee shops, bookstores, enterprises, and restaurants). The experimental results show that 75 (93.75%) out of the 80 evaluated Wi-Fi networks are vulnerable to our TCP hijacking attack.
@@ 2.1 Frame Format and Security Mechanisms in Wi-Fi Network
The 802.11 Frame Layout. Figure 1 shows the layout of the 802.11 frame. Firstly, the Frame Control (FC) field contains several flags and defines the type of the frame. The Type and Subtype fields together identify the function of the frame. There are currently three types (i.e., management, control, and data frames) and more than 50 subtypes defined in 802.11 specifications. In our attack, the attacker needs to monitor the victim’s TCP packets which will be encapsulated into 802.11 frames with type 2 and subtype 8 in Wi-Fi networks. To identify the victim’s encrypted frames, the attacker needs to analyze the addresses of the 802.11 frames. There are four address fields in the 802.11 frame format. These fields are used to indicate the basic service set identifier (BSSID), source address (SA), destination address (DA), transmitting address (TA), and receiving address (RA). Certain frames might not contain some of the address fields. Certain address field usage is specified by the relative position of the address field (1 – 4) within the MAC header, independent of the type of address present in that field. Specifically, the Address 1 field always identifies the intended receiver(s) of the frame, and the Address 2 field, where present, always identifies the transmitter of the frame [3]. In our attack, the attacker can identify the victim supplicant’s encrypted frames through addresses 1 (RA) and address 2 (TA). After filtering the victim’s encrypted frames, the attacker needs to further analyze the payload size of the encrypted frames. The payload (i.e., MSDU in Figure 1) of a normal data frame contains the upper layer data (e.g., TCP packets). The MSDU typically starts with an LLC/SNAP header and is protected by cryptographic encapsulation mechanisms (i.e., TKIP, CCMP, and GCMP). In this paper, we refer to the encrypted frame size as the MSDU size.
@@ 2.2 Challenge ACK Mechanism in TCP
Our attack leverages the trigger conditions of the challenge ACK mechanism in two distinct ways. Firstly, when a receiver detects an incoming SYN packet within an established TCP connection, regardless of the sequence number, it responds by sending an ACK (referred to as the challenge ACK) to the remote peer. This ACK serves as a challenge for the remote peer to confirm the loss of the previous connection and the initiation of a new connection. Only the legitimate peer will receive this ACK and respond with a RST segment containing the correct sequence number, derived from the ACK field of the challenge ACK packet, in the event of connection loss. Consequently, a spoofed SYN packet will generate an additional ACK, which will be disregarded by the peer as a duplicate ACK and will have no impact on the established connection. We will demonstrate how this challenging condition can be exploited to detect a victim TCP connection in Section 4.2.
Secondly, the receiver employs a verification process for the acknowledgment number of each TCP segment to prevent blind data injection attacks. Acceptance of an acknowledgment number for any data segment is contingent upon its falling within the range of ($SND.UNA-SND.WND,SND.NXT$), where $SND.UNA$ represents the sequence number of the first unacknowledged octet, $SND.WND$ denotes the maximum window size observed by the receiver from the sender, and $SND.NXT$ is the next sequence number to be sent. If the acknowledgment number of the segment ($SEG.ACK$) is in the range ($SND.UNA-(2^{31}-1),SND.UNA-SND.WND$), the receiver responds with a challenge ACK. If the $SEG.ACK$ is greater than $SND.NXT$, the receiver silently discards this TCP segment. That can be exploited by attackers to infer the acceptable acknowledgment number, as described in Section 4.3.3.
@@ 2.3 TCP Options
Selective Acknowledgment Option. The TCP selective acknowledgment (SACK) is specified in RFC 2018 [24]. It is an optional feature that is typically enabled by default in the majority of TCP implementations. The SACK option is particularly recommended for networks experiencing frequent packet loss or packet reordering. Its utilization can significantly improve the performance and reliability of TCP connections in such environments. When a receiver detects a TCP segment with a sequence number that has already been acknowledged as outdated, it responds by sending a SACK-ACK to notify the sender. The SACK option will enable the sender to selectively retransmit lost packets based on the information provided by the receiver. However, if the sequence number of the received TCP segment has not yet been acknowledged, the receiver will reply with an ACK packet, which may have a different frame size (as shown in Table 1) or may not respond at all, depending on the acknowledgment number of the segment. In this paper, we will show that attackers in Wi-Fi networks can differentiate between these situations and thus infer the sequence number of a target TCP connection by analyzing the size of encrypted wireless frames.
3 Threat Model
Figure 2 illustrates the threat model of our off-path TCP hijacking attack in Wi-Fi networks. The AP encrypts the network traffic of its supplicants via security mechanisms, e.g., WPA2 or WPA3. A victim supplicant, such as a laptop or a smartphone, connects to the AP and establishes TCP connections with remote servers. The attacker, functioning as a regular supplicant without AP management privileges, utilizes multiple wireless network interface cards (WNICs). One (managed model) of these WNICs connects to the AP, while the others (monitor model) are utilized to sniff encrypted frames transmitted over the shared Wi-Fi channels. We make the assumption that the attacker has prior access to the target Wi-Fi network before performing our attack. This is a commonly accepted assumption in Wi-Fi hijacking scenarios, as highlighted in previous studies [23, 63, 69].
@@ 3 Threat Model
(a) Identifying victim and detecting TCP connections.
(b) Inferring sequence and acknowledgment numbers.
@@ 4.1 Attack Overview
Step 1: Identifying victim. The attacker accesses a Wi-Fi network and scans the WLAN for potential victim supplicants. In this step, the attacker identifies the $<MAC,IP>$ address pair of the victim to monitor its encrypted frames.
@@
Step 3: Inferring Sequence Number. After detecting a victim’s TCP connection, the attacker sends forged TCP packets with guessed sequence numbers to the server. These manipulated TCP packets prompt the server to generate SACK-ACK responses, which will be sniffed by the attacker when they (i.e., 80-byte encrypted frames) are transmitted in the Wi-Fi channel. By monitoring the victim’s encrypted frames, the attacker can identify the correct sequence number of the target TCP connection.
@@
4.2 Identifying victim and Detecting TCP Connections
Identifying victim. The attacker first prepares the TCP hijacking attack from two aspects, i.e., obtaining the $<MAC,IP>$ address pair of the victim and identifying the Wi-Fi channel used by the victim. The attacker actively sends ARP requests in the WLAN to detect other alive supplicants (i.e., the potential victim clients of our TCP hijacking attack). By observing the ARP responses, the attacker can learn the victim’s MAC address and IP address. With the victim’s MAC address, the attacker sniffs encrypted frames in the Wi-Fi channel and filters the victim’s frames based on address 1 (or address 2) in the 802.11 MAC header (see Figure 1). If the Wi-Fi network supports multiple accessed Wi-Fi channels, the attacker scans all Wi-Fi channels to identify the specific channel used by the victim. Subsequently, the attacker intercepts encrypted frames within the target Wi-Fi channel and filters out the victim’s frames.
Detecting TCP Connections. With intercepting and analyzing the victim’s encrypted frames, the attacker can identify the victim’s TCP connections. A TCP connection is recognized by four elements, i.e., [client IP address, client port number, server IP address, server port number]. Typically, the server IP address and server port number are publicly known to the attacker [61, 14, 69]. The attacker needs to infer the client’s IP address and port number. In our attack, the client IP is obtained via ARP response. Thus, the only remaining element to determine is the client port number.
Given that a TCP connection was previously established by the legitimate user on a victim client using a source port p, the attacker impersonates as the client and sends forged SYN/ACK packets to the server. As per the challenge ACK mechanism described in RFC 5961 [59], if the forged SYN/ACK packet contains the same client port number p, the server will respond with a challenge ACK to the client. This challenge ACK packet will be encapsulated into a 68-byte encrypted frame and sniffed by the attacker, during transmission from the AP to the client.
@@ 4.2 Identifying victim and Detecting TCP Connections
The attacker iterates through the above procedure by changing the client port number specified in the forged SYN/ACK packet. This procedure continues until the correct port number p is identified. Finally, the attacker identifies a target TCP connection operating on the four-tuple, i.e., [client IP address, client port number, server IP address, server port number].
@@ 4.3 Inferring Sequence and Acknowledgment Numbers
In this section, we begin with a concise overview of the mechanism used to verify the sequence number and acknowledgment number of TCP segments. Next, we introduce the approach for inferring the exact sequence number and an acceptable acknowledgment number by leveraging encrypted Wi-Fi frames.
@@ 4.3.1 Verifying TCP Segment
According to RFC 9293 [19], upon receiving a TCP segment, the TCP receiver first performs a verification by comparing the sequence number ($SEG.SEQ$) specified in the TCP header with its receive window. In other words, the condition $RCV.NXT\leq SEG.SEQ\leq RCV.NXT+RCV.WND$ must be met, where $RCV.NXT$ denotes the next expected sequence number for an incoming segment, and $RCV.WND$ indicates the size of the receive window. Furthermore, as per the specification, the ACK flag is consistently set to true, except for the initial SYN packet used for connection establishment. If the ACK bit is disabled, the receiver will discard the segment. Therefore, when hijacking the target TCP connection, the attack must infer an acceptable acknowledgment number and sequence number.